Problem # 1: A bond issued on February 1, 2004 with face value of S25400 has semiannual coupons of 7.5%, and can be redeemed for par (face value) on February 1, 2019. What is the accrued interest and the market price (the clean price) of the bond on November 15, 2006, if the bonds yield on that date is to be 9.5%2 (use accrued interest) accrued interest and market price (in that order), separated with a comma both answers correct to 2 decimals Problem #1: Just Save | | Submit Problem #1 for Grading
Answer #1

Face value (F) = redemption value (C) = $25400

Semiannual coupon rate (r) = 0.075/2 = 0.0375

Semiannual yield rate (j) = 0.095/2 = 0.0475

Number of compounding periods till maturity: 30-5 = 25

Number of days since the last coupon = 106

Number of days in a coupon period = 184

Therefore t = 106/184

Therefore the accrued interest is:

tFr = (106/184)(25400)(0.0375) = 548.72

The market price = (price plus accrued at time t) - accrued interest

Price plus accrued at time t = (P(0))((1+j)^t) - tFr

P(0) = F + F(r-j)(PV factor with n=25 and j = 0.0475)

P(0) = 25400 + 25400 (0.0375-0.0475)(14.45) = 21728.69

Therefore the market price is:

21728.69(1.0475^(106/184))-548.72 = 21768.70

answered by: anonymous